# Fermionic Love of Black Holes in General Relativity

### v1

@@ Abstract
In General Relativity, black holes exhibit a remarkable feature: their tidal Love numbers — associated with the conservative response to static tidal fields — vanish identically [3, 4, 5, 6, *Poisson:2014gka, *Landry:2015zfa, 9, 10, 11, *LeTiec:2020spy, 13]. This result stands in sharp contrast with the behavior of other compact objects, which generically have nonzero tidal Love numbers [14]. This notably includes neutron stars – for which tidal Love numbers measured from gravitational-wave mergers [15] play a crucial role in constraining the equation of state [16] – but also other exotic compact objects and black hole mimickers within and beyond General Relativity [17, 18, *Sennett:2017etc, *Mendes:2016vdr, *Uchikata:2016qku, *Raposo:2018rjn, *Cardoso:2019rvt, *Berti:2024moe, *Silvestrini:2025lbe, 14, 26]. Furthermore, the tidal Love numbers are nonzero also for black holes surrounded by matter distributions [27, *DeLuca:2021ite, *DeLuca:2022xlz, *Brito:2023pyl, *Capuano:2024qhv, *Cardoso:2019upw, *Cardoso:2021wlq, *DeLuca:2024uju, *Cannizzaro:2024fpz, *DeLuca:2025bph, 37], for black holes in modified gravity theories and asymptotically nonflat spacetimes [14, 38, *DeLuca:2024nih, *DeLuca:2022tkm, *Barbosa:2025uau, *Barura:2024uog, *nair2024asymptotically-199, *Franzin:2024cah], and finally for higher-dimensional black holes [45, *Chakravarti:2019aup, *Pereniguez:2021xcj, *Dey:2020lhq, *Dey:2020pth, *Cardoso:2019vof, *Rodriguez:2023xjd, *Charalambous:2023jgq, *Charalambous:2024tdj, *Charalambous:2024gpf, *Singha:2025xah, 56, 57, 58].
@@
Indeed, in recent years the vanishing of the black-hole tidal Love numbers has been linked to hidden symmetries of perturbations of the Kerr solution in the zero-frequency limit [57, 60, 61, 62, *BenAchour:2022uqo, 64, *Katagiri:2022vyz, *Ivanov:2022qqt, 67, 68, *Bonelli:2021uvf, *Ivanov:2022hlo, *Rai:2024lho, 13, 72]. These symmetry arguments have also been extended beyond the linear regime [67, 73, 74], and recently shown to persist to all orders in perturbation theory [75, 76, 77, 78] in the static case (while the dynamical tidal response of black holes has been studied in [79, 65, 66, 67, 64, 80, 81, 26, 82, 83, 79, 84, 85, 86, 87] with different motivations and approaches).
@@
Massless perturbations of a Kerr black hole. The unique stationary black hole solution to vacuum General Relativity is the Kerr metric [88, *Robinson], characterized by only two parameters: the mass $M$ and the angular momentum $J=aM$, with $|a|\leq M$. The horizon radii are $r_{\pm}=M\pm\sqrt{M^{2}-a^{2}}$. A remarkable property of this solution is that any perturbation can be decomposed into one or more spin-$s$ scalar wavefunctions $\Psi_{s}$ that admit a separable form,
@@
where $\{t,r,\vartheta,\varphi\}$ are Boyer-Lindquist coordinates, $\omega$ and $m$ denote the frequency and azimuthal number of the perturbation, and $s$ is its spin: integer for bosons and half-integer for fermions. The radial and angular functions, $R_{s}$ and $S_{s}$, are respectively governed by the spin-$s$ Teukolsky equation and by a deformed spin-weighted spherical harmonic equation which, in the massless case, read [90, *Teukolsky:1973ha, 92]
@@
The angular equation for static perturbations reduces to the spin-weighted spherical harmonic equation [94, *1977RSPSA.358...71B], for which regularity at $\cos\vartheta=\pm 1$ requires $|m|\leq\ell$, $\ell\geq|s|$, and the quantization of $(\ell,m)$ in terms of $s$:
@@
A fermionic perturbation: Dirac fields. While the response of the Kerr black hole to massless bosonic probes — scalar ($s=0$), electromagnetic ($s=\pm 1$), and gravitational ($s=\pm 2$) fields — has been extensively studied, one can also consider massless fermionic perturbations, such as Dirac fields ($s=\pm\tfrac{1}{2}$, e.g. neutrinos) [96, *ChandraBook, *Chandrasekhar:1976ap, 99, 100] or Rarita-Schwinger fields ($s=\pm\tfrac{3}{2}$, e.g. gravitinos in supergravity) [101, *TorresdelCastillo:1990aw].
@@
The radial waveforms are directly related to observable quantities: $|\sqrt{\Delta}\,R_{+\frac{1}{2}}|^{2}$ and $|R_{-\frac{1}{2}}|^{2}$ are the probability densities for spin-up and spin-down particles. Moreover, although the Teukolsky equation 2 treats $R_{+\frac{1}{2}}$ and $R_{-\frac{1}{2}}$ as independent, they are in fact related through the underlying first-order system [100]. In the static limit, this implies that the amplitudes $\mathcal{E}_{\pm 1/2\,\ell m}$ are related to each other, even though the corresponding spin-up and spin-down responses $\mathcal{F}_{\pm 1/2\,\ell m}$ may still differ.
@@
In this letter we focused on the static response, but our methods can extend to dynamical settings, generalizing recent analyses [79, 65, 66, 67, 64, 26, 79, 80, *Perry:2023wmm, *Ivanov:2024sds, *DeLuca:2024ufn, *Katagiri:2024wbg, *Katagiri:2024fpn, *Bhatt:2024rpx, *Chakraborty:2025wvs] to the fermionic sector.
Another promising direction is to embed our computation within the worldline effective field theory [93, 107, 59, 57], by analyzing the scattering of massless fermions. This framework would allow one to quantify the imprint of fermionic Love numbers on the dynamics of black-hole binaries. While this scenario requires one of the companions to source a massless fermionic field, it would open the door to novel phenomenology in gravitational-wave observations, potentially connected to recently discovered black-hole solutions with electroweak hair in Einstein–Weinberg–Salam theory [109, 110]. In particular, it would be interesting to investigate the convergence of the post-Newtonian series in the large $\ell=m$ limit, given that the response of near-extremal black holes grows exponentially for both bosonic and fermionic perturbations (see 12 for Dirac fields in the extremal case but this holds also for the bosonic response in 10), suggesting that higher-multipole tidal responses could play an unexpectedly significant role when $a\gtrsim 0.95M$. Such large spins are consistent with estimates for black holes in low-mass X-ray binaries [111, 112], as well as with the spins of the binary components in the recent GW231123 gravitational-wave event [113].
Finally, the vanishing of bosonic Love numbers in General Relativity is also tied to classical no-hair theorems, implying that a black hole cannot sustain static bosonic hair decaying at infinity [114, 115, 61]. Our results highlight the distinctive role of fermions in potentially circumventing these theorems [109, 110], and open new directions to probe the interplay between fundamental fields, black-hole structure, and strong-gravity phenomenology.
Acknowledgments. The research of S.C. is supported by MATRICS (MTR/2023/000049) and Core Research Grants (CRG/2023/000934) from SERB, ANRF, Government of India. P.H. is supported by the Department of Physics at The Ohio State University. P.P. is partially supported by the MUR FIS2 Advanced Grant ET-NOW (CUP: B53C25001080001) and by the INFN TEONGRAV initiative.

### v2

@@ Abstract
In General Relativity, black holes exhibit a remarkable feature: their tidal Love numbers — associated with the conservative response to static tidal fields — vanish identically [1, 11, 31, 42, *Poisson:2014gka, *Landry:2015zfa, 62, 55, 28, *LeTiec:2020spy, 8]. This result stands in sharp contrast with the behavior of other compact objects, which generically have nonzero tidal Love numbers [16]. This notably includes neutron stars – for which tidal Love numbers measured from gravitational-wave mergers [3] play a crucial role in constraining the equation of state [27] – but also other exotic compact objects and black hole mimickers within and beyond General Relativity [74, 63, *Sennett:2017etc, *Mendes:2016vdr, *Uchikata:2016qku, *Raposo:2018rjn, *Cardoso:2019rvt, *Berti:2024moe, *Silvestrini:2025lbe, 16, 20]. Furthermore, the tidal Love numbers are nonzero also for black holes surrounded by matter distributions [4, *DeLuca:2021ite, *DeLuca:2022xlz, *Brito:2023pyl, *Capuano:2024qhv, *Cardoso:2019upw, *Cardoso:2021wlq, *DeLuca:2024uju, *Cannizzaro:2024fpz, *DeLuca:2025bph, 50], for black holes in modified gravity theories and asymptotically nonflat spacetimes [16, 17, *DeLuca:2024nih, *DeLuca:2022tkm, *Barbosa:2025uau, *Barura:2024uog, *nair2024asymptotically-199, *Franzin:2024cah], and finally for higher-dimensional black holes [21, *Chakravarti:2019aup, *Pereniguez:2021xcj, *Dey:2020lhq, *Dey:2020pth, *Cardoso:2019vof, *Rodriguez:2023xjd, *Charalambous:2023jgq, *Charalambous:2024tdj, *Charalambous:2024gpf, *Singha:2025xah, 54, 43, 59].
@@
Indeed, in recent years the vanishing of the black-hole tidal Love numbers has been linked to hidden symmetries of perturbations of the Kerr solution in the zero-frequency limit [43, 24, 44, 7, *BenAchour:2022uqo, 26, *Katagiri:2022vyz, *Ivanov:2022qqt, 33, 25, *Bonelli:2021uvf, *Ivanov:2022hlo, *Rai:2024lho, 8, 69, *Lupsasca:2025pnt]. These symmetry arguments have also been extended beyond the linear regime [33, 67, 45], and recently shown to persist to all orders in perturbation theory [52, 29, 39] in the static case (while the dynamical tidal response of black holes has been studied in [10, 49, 47, 33, 26, 68, 64, 20, 46, 32, 10, 51, 48, 9, 18, 19] with different motivations and approaches).
@@
Massless perturbations of a Kerr black hole. The unique stationary black hole solution to vacuum General Relativity is the Kerr metric [53, *Robinson], characterized by only two parameters: the mass $M$ and the angular momentum $J=aM$, with $|a|\leq M$. The horizon radii are $r_{\pm}=M\pm\sqrt{M^{2}-a^{2}}$. A remarkable property of this solution is that any perturbation can be decomposed into one or more spin-$s$ scalar wavefunctions $\Psi_{s}$ that admit a separable form,
@@
where $\{t,r,\vartheta,\varphi\}$ are Boyer-Lindquist coordinates, $\omega$ and $m$ denote the frequency and azimuthal number of the perturbation, and $s$ is its spin: integer for bosons and half-integer for fermions. The radial and angular functions, $R_{s}$ and $S_{s}$, are respectively governed by the spin-$s$ Teukolsky equation and by a deformed spin-weighted spherical harmonic equation which, in the massless case, read [71, *Teukolsky:1973ha, 70]
@@
The angular equation for static perturbations reduces to the spin-weighted spherical harmonic equation [36, *1977RSPSA.358...71B], for which regularity at $\cos\vartheta=\pm 1$ requires $|m|\leq\ell$, $\ell\geq|s|$, and the quantization of $(\ell,m)$ in terms of $s$:
@@
A fermionic perturbation: Dirac fields. While the response of the Kerr black hole to massless bosonic probes — scalar ($s=0$), electromagnetic ($s=\pm 1$), and gravitational ($s=\pm 2$) fields — has been extensively studied, one can also consider massless fermionic perturbations, such as Dirac fields ($s=\pm\tfrac{1}{2}$, e.g. neutrinos) [14, *ChandraBook, *Chandrasekhar:1976ap, 73, 57] or Rarita-Schwinger fields ($s=\pm\tfrac{3}{2}$, e.g. gravitinos in supergravity) [41, *TorresdelCastillo:1990aw].
@@
The radial waveforms are directly related to observable quantities: $|\sqrt{\Delta}\,R_{+\frac{1}{2}}|^{2}$ and $|R_{-\frac{1}{2}}|^{2}$ are directly related to the probability densities for spin-up and spin-down particles. Moreover, although the Teukolsky equation 2 treats $R_{+\frac{1}{2}}$ and $R_{-\frac{1}{2}}$ as independent, they are related through the first-order system [57]. In the static limit, this implies that the amplitudes $\mathcal{E}_{\pm 1/2\,\ell m}$ are related to each other, even though the corresponding spin-up and spin-down responses $\mathcal{F}_{\pm 1/2\,\ell m}$ may still differ.
@@
Although the fermionic response is extracted from specific spinor components defined with respect to a chosen tetrad, the Love numbers in 10 are invariant under spin-frame transformations. Such transformations act as local Lorentz rotations in spin space. As such, they mix the $R_{\pm 1/2}$ radial solutions linearly but leave unchanged the ratio between the growing and decaying radial modes from which the response coefficients are defined. Hence, the fermionic Love numbers are tetrad- and spin-frame independent. Moreover, parity exchanges the two spin-weight sectors $s\to-s$ (modulo complex conjugation) and leaves the magnitude of the response unchanged, up to the conventional sign associated with spin orientation; so no parity-odd tidal response arises in the static limit, as in the bosonic sector.
Finally, fermionic fields are intrinsically quantum, but bilinears computed from them have a classical counterpart. For $s=1/2$, the simplest bilinear one can construct is $\bar{\Psi}\Psi=\Psi^{\dagger}\gamma^{0}\Psi$. At large distance, using the expressions for $\mathcal{E}_{s\ell m}$ and ${\cal F}_{s\ell m}$, for $s=1/2$ we get
$\bar{\Psi}\Psi\underset{r\to\infty}{\sim}r^{2\ell-1}\Big[$ $(1+{\cal A}+\cdots)$
$+2{\cal A}{\cal F}_{\frac{1}{2}\ell m}\left(\frac{r_{+}}{r}\right)^{2\ell+1}(1+\cdots)\Big]\,$, (13)
where ${\cal A}^{-1}=\left(\frac{1}{4}+4P^{2}\right)\left[\left(\ell+\frac{1}{2}\right)^{2}+4P^{2}\right]$. This shows that the nonvanishing fermionic response affects also classical quantities.
@@
In this letter we focused on the static response, but our methods can extend to dynamical settings, generalizing recent analyses [10, 49, 47, 33, 26, 20, 10, 68, *Perry:2023wmm, *Ivanov:2024sds, *DeLuca:2024ufn, *Katagiri:2024wbg, *Katagiri:2024fpn, *Bhatt:2024rpx, *Caron-Huot:2025tlq, *Chakraborty:2025wvs] to the fermionic sector.
Another promising direction is to embed our computation within the worldline effective field theory [38, 37, 66, 43], by analyzing the scattering of massless fermions.
This framework would allow one to quantify the imprint of fermionic Love numbers on the dynamics of black-hole binaries. While this scenario requires one of the companions to source a massless fermionic field, it would open the door to novel phenomenology in gravitational-wave observations, potentially connected to recently discovered black-hole solutions with electroweak hair in Einstein–Weinberg–Salam theory [35, *Gervalle:2025awa]. In particular, it would be interesting to investigate the convergence of the post-Newtonian series in the large $\ell=m$ limit, given that the response of near-extremal black holes grows exponentially for both bosonic and fermionic perturbations (see 12 for Dirac fields in the extremal case but this holds also for the bosonic response in 10), suggesting that higher-multipole tidal responses could play an unexpectedly significant role when $a\gtrsim 0.95M$. Such large spins are consistent with estimates for black holes in low-mass X-ray binaries [12, 60], as well as with the spins of the binary components in the recent GW231123 gravitational-wave event [2].
Finally, the vanishing of bosonic Love numbers in General Relativity is also tied to classical no-hair theorems, implying that a black hole cannot sustain static bosonic hair decaying at infinity [5, *Bekenstein:1972ky, 44]. Our results highlight the distinctive role of fermions in potentially circumventing these theorems [35, *Gervalle:2025awa], and open new directions to probe the interplay between fundamental fields, black-hole structure, and strong-gravity phenomenology. For example, the nonzero fermionic response indicates that the lowest mode at $\ell=|s|$ corresponds to a static, normalizable mode (see 4), i.e. a fermionic hair. In particular, for spin-$\tfrac{1}{2}$, we find that the four-spinor does not diverge at infinity while remaining nontrivial at the horizon for $\ell=\tfrac{1}{2}$. The existence of such a mode could have been anticipated from supersymmetric theories of gravity: since General Relativity can be embedded into four-dimensional supergravity, the fermionic superpartners of the bosonic (spin-$1$ and spin-$2$) fields defining the Kerr–Newman solution guarantee the existence of a spin-$\frac{1}{2}$ and spin-$\frac{3}{2}$ normalizable perturbation at the lowest value of $\ell$ in Kerr–Newman. An open question raised by our work is whether nontrivial fermionic hair can also be constructed for the full tower of $\ell$ values for which the response is nonzero, and whether they survive at the nonlinear level.
Acknowledgments. We thank Miguel Correia, Gerardo Garcia Moreno, and Samir Mathur for interesting comments. The research of S.C. is supported by MATRICS (MTR/2023/000049) and Core Research Grants (CRG/2023/000934) from SERB, ANRF, Government of India. P.H. is supported by the Department of Physics at The Ohio State University. P.P. is partially supported by the MUR FIS2 Advanced Grant ET-NOW (CUP: B53C25001080001) and by the INFN TEONGRAV initiative.